Calculate the kinetic energy of an object of mass 500g that is moving with a velocity of 20m/s
1
Expert's answer
2020-05-14T09:11:28-0400

The kinetic energy


"KE=\frac{mv^2}{2}=\frac{0.5\:\rm kg\times (20\: m/s)^2}{2}=100\:\rm J"